OpenStudy (anonymous):

find the 5th term in the expansion of (3x^2-sqrt(y))^6 a)135x^4y^2 b)45x^4y^2 c)-135x^4y^2 d)-45x^4y^2

OpenStudy (alexwee123):

binomial theorem anyone?

OpenStudy (anonymous):

i think its a)135x^4y^2 right?

OpenStudy (dumbcow):

exponent is 6 term number is 5 \[T_{5} = \left(\begin{matrix}6 \\ 5\end{matrix}\right)a^{6-5} b^{5}\] where a = 3x^2 b = -sqrt(y)

OpenStudy (dumbcow):

wait i did something wrong there the answer is A

OpenStudy (dumbcow):

it should be \[\left(\begin{matrix}6 \\ 4\end{matrix}\right)(3x^{2})^{6-4} (-\sqrt{y})^{4} = 15*9*x^{4}*y^{2}\]